# Transmission Coefficient of Super-Janus Solution Saba Baig, Andreas Karch, Mianqi Wang

### v1

@@ 2 CFT calculation
More generally, one can obtain a class of ICFTs by adding in an exactly marginal deformation, which in turn gives the gluing condition at the boundary/interface [18]. In addition to the radii deformation above, let us consider a $\mathbb{Z}_{2}$ orbifold twist deformation. It is obtained by perturbing the theory with a dimension $(h,\bar{h})=(1,1)$ twist operator $T_{0}$, with an OPE of $T_{0}(z,\bar{z})T_{0}(w,\bar{w})\sim 1/|z-w|^{4}$. Under this twist, the orbifold CFTs on the two sides live on different points in their moduli space. Following [18] it was further argued that the orbifold twist deformation is one of the many possible Kähler moduli deformations. By counting distinct two forms on the orbifold point of a Kähler manifold (e.g. a $T^{4}/\mathbb{Z}_{2}$ orbifold point in $K_{3}$), one can naturally extrapolate the presence of orbifold twist-fields to jumping world-sheet $B$ fields across the interface.
@@
$\left(\begin{array}[]{c}\partial_{-}X^{1}\\ \partial_{+}\tilde{X}^{1}\\ \partial_{-}X^{2}\\ \partial_{+}\tilde{X}^{2}\\ \end{array}\right)=S\left(\begin{array}[]{c}\partial_{+}X^{1}\\ \partial_{-}\tilde{X}^{1}\\ \partial_{+}X^{2}\\ \partial_{-}\tilde{X}^{2}\\ \end{array}\right)$ (2.7)
where we only write out the $2\times 2$ transmission submatrix for $\partial_{-}X^{i}$ and $\partial_{-}\tilde{X}^{i}$
@@
Following [9], we can fold the theory via $\sigma\to-\sigma$ (and adding a minus sign in front of the $B$ field), and calculate the energy transmission coefficient using the above S matrix on the conformal boundary state $|b\rangle$. Explicitly, if we set the boundary condition to be
$(L_{m}^{i}-S_{ij}\bar{L}_{-m}^{j})|b\rangle=0,$ (2.9)
then the energy transmission coefficient is
$\mathcal{T}_{\text{weak}}=\frac{2}{c_{1}+c_{2}}\left(\frac{\langle 0|L_{2}^{1}\bar{L}_{2}^{2}|b\rangle}{\langle 0|b\rangle}+\frac{\langle 0|L_{2}^{2}\bar{L}_{2}^{1}|b\rangle}{\langle 0|b\rangle}+\frac{\langle 0|L_{2}^{1}\bar{L}_{2}^{4}|b\rangle}{\langle 0|b\rangle}+\frac{\langle 0|L_{2}^{4}\bar{L}_{2}^{1}|b\rangle}{\langle 0|b\rangle}\right)$ (2.10)
Plugging in the S matrix leads to the result

### v2

@@ 2 CFT calculation
More generally, one can obtain a class of ICFTs by adding in an exactly marginal deformation, which in turn gives the gluing condition at the boundary/interface [18]. In addition to the radii deformation above, let us consider a $\mathbb{Z}_{2}$ orbifold twist deformation [23]. It is obtained by perturbing the theory with a dimension $(h,\bar{h})=(1,1)$ twist operator $T_{0}$, with an OPE of $T_{0}(z,\bar{z})T_{0}(w,\bar{w})\sim 1/|z-w|^{4}$. Under this twist, the orbifold CFTs on the two sides live on different points in their moduli space. Following [18] it was further argued that the orbifold twist deformation is one of the many possible Kähler moduli deformations. By counting distinct two forms on the orbifold point of a Kähler manifold (e.g. a $T^{4}/\mathbb{Z}_{2}$ orbifold point in $K_{3}$), one can naturally extrapolate the presence of orbifold twist-fields to jumping world-sheet $B$ fields across the interface.
@@
$\left(\begin{array}[]{c}\partial_{-}X^{1}\\ \partial_{-}X^{2}\\ \partial_{+}\tilde{X}^{1}\\ \partial_{+}\tilde{X}^{2}\\ \end{array}\right)=S\left(\begin{array}[]{c}\partial_{+}X^{1}\\ \partial_{+}X^{2}\\ \partial_{-}\tilde{X}^{1}\\ \partial_{-}\tilde{X}^{2}\\ \end{array}\right)$ (2.7)
where we only write out the $2\times 2$ transmission submatrix $S_{21}$ for $\partial_{-}X^{i}$ and $\partial_{-}\tilde{X}^{i}$
@@
Following [9], we can fold the theory via $\sigma\to-\sigma$ (and adding a minus sign in front of the $B$ field), and calculate the energy transmission coefficient using the above S matrix on the conformal boundary state $|b\rangle$. Explicitly, we can write down the $2\times 2$ $R$-matrix
$R_{ij}=\frac{\langle 0|L_{2}^{i}\bar{L}_{2}^{j}|b\rangle}{\langle 0|b\rangle},$ (2.9)
where $L_{2}^{1,2}$ are the virasoro generators for the left and right hand side. The energy transmission coefficient is then
$\mathcal{T}_{\text{weak}}=\frac{2}{c_{1}+c_{2}}\left(\frac{\langle 0|L_{2}^{1}\bar{L}_{2}^{2}|b\rangle}{\langle 0|b\rangle}+\frac{\langle 0|L_{2}^{2}\bar{L}_{2}^{1}|b\rangle}{\langle 0|b\rangle}\right)=\frac{1}{2}\sum_{I,J=1,2}\big{|}(S_{21})_{IJ}\big{|}^{2}$ (2.10)
Plugging in (2.8) leads to the result